# Towards Generalist Robot Policies: What Matters in Building Vision-Language-Action Models

### v3

@@ Appendix C Implementation Details
With different formulations, the best setting of hyperparameters like batch size, weight decay, and learning rate could be varied. Although OpenVLA suggests utilizing the same hyper-param as in the VLM pretrain phase, we find that a varied setting of the hyper-param could improve the performance.
The hyperparameters for fine-tuning VLAs are mainly derived from the VLMs training setups, for example, we select the weight decay from , and the learning rate as one of . We conduct a grid search over and select the one with the best performance. We set the global batch size as 128 and the warmup ratio is 0.25 epoch (5K steps for OpenX Embodiment pre-train). All models included in this paper are trained on a cluster of 4 x 8 A100 GPUs.
@@
| CALVIN Perform (Tab. II) | All | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| SimplerEnv Perform (Fig. 14) | All | 16 | 10 | Side | 128 | 5K Iters | Constant | AdamW | 1e-4 | 50K Iters |
| CALVIN VL Pre-train (Fig. 6) | All | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| Real Perform (Fig. 7) | All | 8 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| VLA Structure (Tab.III) | LLaVA | 8 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 2e-5 | 5 Ep |
| Else | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep | |
| CALVIN Generalization (Fig. 9) | All | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| CALVIN Data Efficiency (Tab. IV) | All | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| CALVIN Backbone (Tab V) | All | 8 | 10 | Side | 128 | 0.25 Ep | Constant | AdamW | 2e-5 | 5 Ep |
| Simpler Training Recipe (Fig 10) | All | 16 | 10 | Side | 128 | 5K Iters | Constant | AdamW | 2e-5 | 50K Iters |
| CALVIN few-shot (Fig. 11) | All | 16 | 10 | Side | 128 | 0 Iter | Constant | AdamW | 2e-5 | 5K Iters |
@@
We find out that, normally, the performance of robot policies does not fully depend on offline evaluation metrics [16], such as the validation loss, due to the compounding error in long-horizon rollouts. Therefore, it is challenging to select the best checkpoint during training. For fair comparisons, we train all VLAs for a fixed number of epochs or timesteps. Concretely, on CALVIN, we train each model for 5 epochs with a batch size of 128 truncated trajectories and report the performance of the final model. For SimplerEnv, we train the model for 100K iterations with a batch size of 512 truncated trajectories and report the best-performing model with a 10K-iteration training interval. In real-world experiments, we train the model for 5 epochs with a batch size of 512 truncated trajectories, and we only report the performance of the last model.
Appendix D Benchmark Details
CALVIN [32] is a simulated benchmark for multi-task table-top manipulation. It provides 24k human-teleoperated demonstrations annotated with language instruction. Each trajectory is less than 64-time steps, which includes 34 pre-defined basic skills: rotate blue block right, move slider right, lift red block slider, place slider, turn off light bulb, turn off led light, push in drawer, lift blue block drawer, close drawer, lift pink block slider, lift pink block table, move slider left, open drawer, turn on light bulb, rotate blue block left, push blue block left,rotate red block right, turn on led light,
@@ Appendix D Benchmark Details
The metrics are the the success rates of finishing these sequential tasks and the average length of achieved tasks.
@@
SimplerEnv [25] are designed as a suite of real-to-sim environments, which enables evaluating robotic policies in simulation as efficient, scalable, and informative alternative to real-world evaluation. SimplerEnv created a comparable arena for benchmarking robotics policies on private real-world setups as Google [6, 7] and BridgeData V2 [45].
@@
Appendix E Detailed Performance on CALVIN
@@
Appendix F Diverse Backbone
@@ Appendix G Diverse ph
Appendix H Detailed Performance on SimplerEnv
| RT-1-X | 0.167 | 0.000 | 0.208 | 0.042 | 0.083 | 0.000 | 0.000 | 0.000 |
| Octo-Base | 0.347 | 0.125 | 0.528 | 0.083 | 0.319 | 0.000 | 0.667 | 0.431 |
| Octo-Small | 0.778 | 0.472 | 0.278 | 0.097 | 0.403 | 0.042 | 0.875 | 0.569 |
| OpenVLA-7b | 0.041 | 0.000 | 0.333 | 0.000 | 0.125 | 0.000 | 0.083 | 0.041 |
| RoboVLMs (Ours) | 0.708 | 0.458 | 0.333 | 0.208 | 0.542 | 0.042 | 0.917 | 0.792 |
Standing | Average | Average | Open | Close | Average | Average | ||||
| RoboVLMs (Ours) | 0.940 | 0.470 | 0.910 | 0.773 | 0.617 | 0.333 | 0.531 | 0.435 | 0.241 | |||
Appendix I Sub-task Performance with Cross-Embodiment Dataset
| Cross-Emb Pre-Train | 0.375 | 0.208 | 0.333 | 0.250 | 0.083 | 0.083 | 0.000 | 0.000 |
| In-domain Full Finetune | 0.542 | 0.292 | 0.250 | 0.250 | 0.458 | 0.125 | 0.583 | 0.583 |
| Post Train | 0.708 | 0.458 | 0.333 | 0.208 | 0.542 | 0.042 | 0.917 | 0.792 |
Standing | Average | Average | Open | Close | Average | Average | ||||
| Cross-Emb Pre-Train | 0.850 | 0.430 | 0.900 | 0.727 | 0.663 | 0.287 | 0.25 | 0.268 | 0.361 | |||
| Target Task Finetune | 0.210 | 0.160 | 0.260 | 0.210 | 0.292 | 0.167 | 0.667 | 0.417 | 0.000 | |||
| In-domain Full Finetune | 0.920 | 0.810 | 0.980 | 0.903 | 0.625 | 0.213 | 0.454 | 0.333 | 0.056 | |||
| Post Train | 0.940 | 0.470 | 0.910 | 0.773 | 0.617 | 0.333 | 0.531 | 0.435 | 0.241 | |||
Appendix J Rollout Examples in SimplerEnv
Appendix K Rollout Examples in Real-World Experiements

### v4

@@ Appendix B Implementation Details
With different formulations, the best setting of hyperparameters like batch size, weight decay, and learning rate could be varied. Although OpenVLA suggests utilizing the same hyperparameters as in the VLM pretrain phase, we find that a varied setting of the hyperparameters could improve the performance.
The hyperparameters for fine-tuning VLAs are mainly derived from the VLMs training setups, for example, we select the weight decay from , and the learning rate as one of . We conduct a grid search and select the one with the best performance. We set the global batch size as 128 and the warmup ratio is 0.25 epoch (5K steps for OpenX Embodiment pre-train). All models included in this paper are trained on a cluster of 4 x 8 A100-80G GPUs.
@@
| CALVIN Perform (Extended Table. 2) | KosMos-2 | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| SimplerEnv Training Recipe (Supplementary Fig. 1, Extended Fig.4) | PaliGemma (Co-Train) | 1 | 10 | Side | 1024 | 500 Iters | Constant | AdamW | 2e-5 | 20K Iters |
| PaliGemma (Post-Train & Finetune) | 1 | 10 | Side | 128 | 0.25 Ep | Constant | AdamW | 5e-5 | 5 Ep | |
| CALVIN VL Pre-train (Fig. 4b) | Flamingo & KosMos-2 | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| Real Perform (Fig. 4d) | KosMos-2 | 8 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| VLA Structure (Table. I) | LLaVA | 8 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 2e-5 | 5 Ep |
| Flamingo & KosMos-2 & PaliGemma | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep | |
| CALVIN Generalization (Fig. 4c) | All | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| CALVIN Data Efficiency (Extended Table. 4) | All | 16 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 1e-4 | 5 Ep |
| CALVIN Backbone (Extended Table. 5) | All | 8 | 10 | Side | 128 | 0.25 Ep | Constant | AdamW | 2e-5 | 5 Ep |
| CALVIN Training Objective (Table. IIa) | All | 1 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 2e-5 | 5 Ep |
| CALVIN MoE (Table. IIb) | All | 1 | 10 | Side+Wrist | 128 | 0.25 Ep | Constant | AdamW | 2e-5 | 5 Ep |
| CALVIN few-shot (Extended Fig. 5) | All | 16 | 10 | Side | 128 | 0 Iter | Constant | AdamW | 2e-5 | 5K Iters |
@@
We find out that, normally, the performance of robot policies does not fully depend on offline evaluation metrics, such as the validation loss, due to the compounding error in long-horizon rollouts. Therefore, it is challenging to select the best checkpoint during training. For fair comparisons, we train all VLAs for a fixed number of epochs or timesteps. Concretely, on CALVIN, we train each model for 5 epochs with a batch size of 128 truncated trajectories and report the performance of the final model. For SimplerEnv, we train the model for 100K iterations with a batch size of 512 truncated trajectories and report the best-performing model with a 10K-iteration training interval. In real-world experiments, we train the model for 5 epochs with a batch size of 512 truncated trajectories, and we only report the performance of the last model.
Appendix C Benchmark Details
CALVIN is a simulated benchmark for multi-task table-top manipulation. It provides 24k human-teleoperated demonstrations annotated with language instruction. Each trajectory is less than 64-time steps, which includes 34 pre-defined basic skills: rotate blue block right, move slider right, lift red block slider, place slider, turn off light bulb, turn off led light, push in drawer, lift blue block drawer, close drawer, lift pink block slider, lift pink block table, move slider left, open drawer, turn on light bulb, rotate blue block left, push blue block left,rotate red block right, turn on led light,
@@ Appendix C Benchmark Details
The metrics are the success rates of finishing these sequential tasks and the average length of completed tasks.
@@
SimplerEnv are designed as a suite of real-to-sim environments, which enables evaluating robotic policies in simulation as efficient, scalable, and informative alternative to real-world evaluation. SimplerEnv created a comparable arena for benchmarking robotics policies on private real-world setups as Google and BridgeData V2.
@@
Appendix D Detailed Performance on CALVIN
@@
Appendix E Diverse Backbone
@@ Appendix F Diverse ph
Appendix G Rollout Examples in SimplerEnv
Appendix H Rollout Examples in Real-World Experiments